When you try to stop your car in an emergency, there is some time before you can react. Your car is going 25 m/s and your reaction time is 0.20 s, then after you hit your brakes it decelerates at 9.5 m/s2. How far will your car travel before it stops?
1
Expert's answer
2020-10-07T07:25:48-0400

During the first 0.2s, before your reacted, the car had passed


"d_1 = 25m/s\times 0.2s = 5m"

After you hit the brakes it decelerates in the following time before the full stop:


"v_0 = at\\
t = \dfrac{v_0}{a}"

where "v_0 = 25m\/s" and "a = 9.5m\/s^2".

In this time the car passes the following distance:

"d_2 = v_0t - \dfrac{at^2}{2} = \dfrac{v_0^2}{a} - \dfrac{v_0^2}{2a} = \dfrac{v_0^2}{2a}"

Obtain:


"d_2 = \dfrac{v_0^2}{2a} = \dfrac{25^2}{2\cdot 9.5} \approx 32.9m"

The full distance:


"d = d_1 + d_2 = 5+32.9 = 37.9m"

Answer. 37.9 m.